Amy is 5 years younger than bob. in three years bob will be twice Amy's age. Use x as bob's age. Write an equation to figure out Amy and Bob's ages
Answer by edjones(8007)

a=b-5
a+3=2b
.
b-5+3=2b Substitute b-5 for a. You can't write 1 equation to find out both ages because you have 2 variables. With this equation you find out bob's age.
.
Ed